### v2

# SpecGen: Automated Generation of Formal Program Specifications via Large Language Models

Abstract
In the software development process, formal program specifications play a crucial role in various stages, including requirement analysis, software testing, and verification. However, manually crafting formal program specifications is rather difficult, making the job time-consuming and labor-intensive. Moreover, it is even more challenging to write specifications that correctly and comprehensively describe the semantics of complex programs. To reduce the burden on software developers, automated specification generation methods have emerged. However, existing methods usually rely on predefined templates or grammar, making them struggle to accurately describe the behavior and functionality of complex real-world programs.
To tackle this challenge, we introduce SpecGen, a novel technique for formal program specification generation based on Large Language Models (LLMs). Our key insight is to overcome the limitations of existing methods by leveraging the code comprehension capability of LLMs. The process of SpecGen consists of two phases. The first phase employs a conversational approach that guides the LLM to generate appropriate specifications for a given program, aiming to utilize the ability of LLM to generate high-quality specifications. The second phase, designed for where the LLM fails to generate correct specifications, applies four mutation operators to the model-generated specifications and selects verifiable specifications from the mutated ones through a novel heuristic selection strategy by assigning different weights of variants in an efficient manner. We evaluate SpecGen on two datasets, including the SV-COMP Java category benchmark and a manually constructed dataset containing 120 programs. Experimental results demonstrate that SpecGen succeeds in generating verifiable specifications for 279 out of 385 programs, outperforming the existing purely LLM-based approaches and conventional specification generation tools like Houdini and Daikon. Further investigations on the quality of generated specifications indicate that SpecGen can comprehensively articulate the behaviors of the input program.
I Introduction
Formal specifications play a central role in describing, understanding, and reasoning about program behaviors. They capture the intended or actual program behaviors, in terms of formal languages, with precise semantics. Formal specifications may take various forms, such as procedure pre-/post-conditions, loop invariants, and assertions at specific program locations. They are essential in a variety of software quality assurance tasks, including software testing [1, 2], model checking [3, 4], and program verification [5, 6].
Yet, a practical challenge is the absence of documented formal specifications in most real-world software projects, since manually writing high-quality specifications is highly nontrivial. To alleviate the burden on software developers, several tools have been introduced for generating program specifications automatically [7, 8, 9], including Houdini [7] and Daikon [9], two most representative ones for Java programs. However, these tools rely heavily on predefined templates or grammars during the specification generation process. This often results in overly simplistic specifications that struggle to accurately capture the complex behaviors and functionalities of real-world programs. Consequently, the specifications generated by such tools usually lack practical significance in the actual software development process.
To address this challenge, we introduce SpecGen, an automated technique for Java program specification generation based on the Large Language Models (LLMs). With the rise of LLMs, extensive research has attempted to apply them in software engineering and LLMs exhibit outstanding performance in various tasks [10, 11, 12, 13, 14], where LLMs have demonstrated remarkable capabilities on code comprehension and summarization [15]. Inspired by this insight, we believe that LLMs can serve as a potent solution to overcome the limitations of existing program specification generation methods. The core idea of this work is to leverage LLMs to generate specifications that accurately capture the real behaviors of input programs, thus imbuing these specifications with richer semantics for further practical use.
The workflow of SpecGen comes in two phases. In the first phase, conversation-driven specification generation, we aim to query the output specifications by conducting a conversation with the LLM. To start the conversation, a prompt is constructed with several few-shot examples for the initial query. During the conversation process, we utilize the verification failure information from the specification verifier as the feedback prompt for the next round of the conversation. In this way, LLMs receive more cues, facilitating them to better generate accurate specifications. Nevertheless, despite the powerful code understanding and generation capabilities of large language models, they still struggle to handle complex programs effectively i.e., generating accurate specifications for complex programs. Through our repeated observation and testing of the model-generated results, we found that although the generated content is not highly accurate, it is already very close to the oracle, which motivates us to design the second phase, mutation-based specification generation. It focuses on generating accurate specifications where the LLM fails to provide verifiable results. Specifically, given a verification failure result by the LLM, SpecGen endeavors to combine four different kinds of mutation operators to modify it and obtain all potential variants. A selector adopting a heuristic selection strategy by assigning different weights of variants further repeatedly chooses a subset of these mutated variants deemed most likely to pass the verification until the results are successfully verified.
To evaluate SpecGen, we conduct experiments on two datasets. We first evaluate SpecGen on the benchmark for the Java category of SV-COMP [16]. To further evaluate the performance of SpecGen on different kinds of programs, we constructed another dataset containing 120 Java programs with manually written ground-truth specifications. The selected programs are highly representative, encompassing different control-flow structures and various data structures to avoid any bias in our evaluation. We compared the performance of SpecGen on the dataset against three baselines: Houdini, Daikon, and Purely LLM-based methods. The results of our evaluation demonstrate that SpecGen significantly outperforms the baseline methods. SpecGen successfully generated verifiable specifications for 279 out of the total 385 programs, outweighing 218 for conversational specification generation, the best-performing Purely LLM-based method, and 98 for Houdini, the best-performing non-LLM method. An ablation study on mutations was also conducted, proving the effectiveness of all four types of mutation operators. Additionally, the results of evaluations on the heuristic selection strategy suggested that our strategy effectively improves the efficiency of SpecGen compared to the random selection strategy. Furthermore, a user study was conducted to evaluate the semantic quality of the generated specifications, illustrating the ability of SpecGen to accurately and comprehensively characterize program behaviors. The main contributions are summarized as follows:
  • A novel approach for formal program specification generation leveraging the Large Language Models to generate accurate and comprehensive specifications to describe program behaviors. Benefiting from the code comprehension ability of LLMs, our approach is capable of generating specifications with high quality, overcoming the limitations of existing methods in generating simplistic and basic specifications.
  • A mutation-based generation approach to enrich the diversity of the LLM output, consisting of a set of mutation operators and a novel heuristic selection strategy proposed to improve the efficiency of the verification that existing works fail to consider.
  • A dataset namely SpecGenBench with hand-written specifications by experts, facilitating follow-up research. Other than the established benchmark SV-COMP, we collected programs on a more diverse spectrum for deeper insights.
  • A comprehensive evaluation from different aspects to evaluate the effectiveness of our approach. We compare SpecGen against Purely LLM-based approaches and representative non-LLM approaches. SpecGen succeeds in 279 out of the 385 programs, significantly outperforming the baseline approaches. Our code and data are available at [17].
II Background and Motivation
II-A Specification Generation and Verification
Program specifications encompass precise statements that describe the intended or actual behaviors of a particular program, either in its entirety or in distinct parts. In this work, we focus on generating specifications for the actual behaviors of input programs. A large proportion of program specifications are expressed in formal languages, such as mathematical expressions to describe the constraints on the behaviors of a program. There are different kinds of specifications such as pre-conditions, which establish constraints on function parameters, ensuring proper execution of the function, post-conditions, which delineate the properties of a set of variables that persist after a function is executed, and loop invariants, which represent a specialized form of specification, detailing properties that consistently hold before executing the loop body. For different programming languages, the specifications may have different implementation forms. For example, in Java, the specifications can be expressed in Java Modeling Language (i.e., JML) [18] where requires statements denote the pre-conditions of a function, ensures statements represent the post-conditions of a function, and maintaining statements are the loop invariants in the program.
[…]
II-B Motivation
Figure 1: An example program and corresponding specifications generated by SpecGen, for which existing tools cannot generate comprehensive specifications to describe the program behaviors.
[…]
Recently, large language models (i.e., LLMs) [23, 24] have exhibited powerful capacities in coding [15, 14]. The emergence of these models may greatly compensate for the limitations of traditional software analysis tools in code understanding. A significant amount of work attempts to leverage large language models in software engineering [25, 26, 27, 28, 29] and we have witnessed substantial progress brought about by the introduction of LLMs. Inspired by these works, in this work, we aim to leverage the large language models in the automated generation of formal program specifications to address the limitations of conventional templated-based approaches. From this perspective, we innovate our approach SpecGen, which generated three parts of specifications for the example presented in the right part of Figure 1. The first part is to describe method TwoSum() as a whole, specifying its pre-conditions and post-conditions. The specifications in lines a and b claim that the input array must not be null before and after the method is executed. The post-condition in line c specifies that the target value equals the sum of the two elements corresponding to the indexes stored in the returned array. The post-condition at line d specifies that there does not exist such a pair of elements that satisfies the constraint when the length of the returned array is zero. The functionality of method TwoSum() can be fully articulated by these generated specifications. Furthermore, the loop invariants in the second and third parts specify corresponding constraints that must be met within a certain range in the array. These generated specifications by SpecGen comprehensively describe the semantics of this function and their correctness is verified by the verifier.
III Approach
III-A Overview
Figure 2: Overview of our SpecGen.
Figure 2: Overview of our SpecGen.
Figure 3: Conversational generation.
The overview of SpecGen is presented in Figure 3, which consists of two components i.e., conversation-driven specification generation and mutation-based specification generation. For the former, SpecGen is designed to communicate with the large language model to query the output in a conversational manner. In particular, a prompt is constructed with some few-shot examples for the initial query. The verification failure information provided by the verifier is further used as the prompt for the next round of conversation if the model-generated results are incorrect. The conversation will be repeated iteratively until the generated specifications successfully pass the verifier or a maximum number of iterations is reached. For the latter, which aims at generating the specifications of a program that the large language model fails to generate. Four kinds of mutation operators are adopted to mutate the specification that failed verification by the verifier and obtain all potential variants. A heuristic selector is further designed to efficiently choose a set of mutated variants that is most likely to pass verification.
III-B Conversation-Driven Specification Generation
Engaging in conversation with large models can fully leverage their capabilities, better assisting them in generating the desired content and avoiding potential errors [30]. Inspired by Xia et al. [11], we propose our conversation-driven specification generation in SpecGen to interact with the LLM conversationally to generate specifications. There are two main benefits: firstly, the conversational manner aids the large model in automatically correcting potential syntax errors in the generated content, secondly, providing the model with the verification failures information by the verifier in the conversation helps it generate more accurate specifications. The design mainly consists of two sequential components: initial prompt construction, which pre-defined an initial prompt to prepare for querying with the LLM, and conversational specification generation, which communicates to the LLM by incorporating verification failure information produced by the verifier in the conversation manner to generate verifiable specifications. The conversation will be repeated iteratively until the generated specifications pass the verifier or a maximum number of iterations is reached.
Figure 4: Illustration of the initial prompt construction.
III-B1 Initial Prompt Construction
We need to define the initial prompt to query with LLM to obtain the model output. After multiple attempts to assess the impact of different prompts on the quality of generated program specifications, we ultimately chose to follow Xia et al. [31] in designing our prompt. The prompt is presented in Fig. 4 illustrating the components of the initial prompt, which consists of three different parts: the system role, few-shot examples, and the queried program. The system’s role aims to inform LLMs about our application scenario, which is to generate JML specifications. We further add some few-shot examples. The reasons are two-fold. On one hand, few-shot examples can help the model to generate more accurate outputs [32]. On the other hand, LLMs can generate the desired output format that is learned from these examples. Each example is a pair of a program and its corresponding specifications. We randomly select it from our collected dataset to construct the few-shot examples. The last component is the queried program which requires the model to generate the output.
III-B2 Conversational Specification Generation
Given the initial prompt, LLM can obtain the initial output for the input program. As the output of the LLM in the first attempt may not successfully pass the validation of the verifier, we interleave the process of specification generation with verification failure feedback to prompt future generation in a conversational manner which is illustrated in Fig. 3. In particular, each generated specification by the model is verified by a JML verifier to test whether the generated result can pass the verifier. If the verification fails, we construct feedback information using the reported error message from the verifier as the prompt for the next generation. The verification error message can help the model understand the reason for failure and provide guidance for generating correct specifications. In addition, to avoid the verifier providing excessively long error messages, we configure the verifier to report only one verification failure message per attempt. Furthermore, through a massive amount of experiments, we summarize several types of common verification failures reported by the verifier. For each kind of error, we provide guidance in the natural language to facilitate the model in generating correct specifications. Detailed guidance information is provided on our website [17]. Upon encountering these types of verification failures reported by the verifier, we will insert corresponding guidance information into the prompt e.g., $<$Guidance$>$ in Fig. 3 to assist the model in resolving the issues. The conversation will be repeated iteratively until the specifications are successfully verified or a maximum number of iterations is reached.
III-C Mutation-based Specification Generation
In the conversation-driven generation process, some few-shot examples are provided in the initial prompt to start the query with the LLM. To further stimulate the potential of the LLM, multi-turn conversation continually guides the LLM to approach the accurate specification more closely. Yet, they still struggle to generate fully correct specifications for some complex programs. The reasons are two-fold. On one hand, in comparison to code generation [32, 33, 34], specification generation poses greater challenges, even for LLMs, because there are currently limited corpora related to program specifications for the model to learn from. On the other hand, although the verification failure information provided by the verifier can assist LLMs in providing higher-quality responses to some extent, as the error messages are highly abstract and generalized, LLMs still struggle to accurately understand the semantic information within error messages for complex programs, leading to incorrect specifications. While LLMs may not accurately generate specifications for complex programs, we found that the generated results are already close enough to the oracle, inspiring us to design the mutation-based generation component.
In particular, the mutation-based specification generation component takes the output generated by the large language model that fails to pass the verifier through the multi-round conversation as the input. We further define a set of mutation operators to modify these generated outputs to obtain more diverse results. Then a heuristic strategy is adopted for efficient verification. The workflow is presented in Algorithm 1. Specifically, we define the specifications generated by LLM that fail verification as the set of template specifications $E_{t}$, which consists of different specifications generated for different locations in a program, and a set of mutation operators as $M$. The MutationBasedGen takes $E_{t}$ and $M$ as the input and outputs a set of correct specifications as $E$. The function SpecMutation corresponds to the mutation operation of $E_{t}$, where each kind of mutation operator will be performed through the mutation function $Mutate()$ (Section III-C1) on a template specification $e$ ($e\in E_{t}$) to obtain a set of candidates $E_{mutated}$ (lines 7 and line 8).
After the mutation operations are performed, we further design the specification selection algorithm to select a subset $E_{selected}$ of mutated specifications that can pass the verification. The selected subset $E_{selected}$ is initialized with $E_{t}$. We then iteratively require the verifier to check the correctness of $E_{selected}$ and obtain a set of refuted specifications denoted as $E_{refuted}$ from $E_{selected}$ that the verifier fails to verify. After that, we need to replace the failed specifications from $E_{refuted}$ with another mutated variant for the next iteration of verification. The $ReSelect$ function is presented from line 18 to line 25. For each refuted specification $e_{r}\in E_{refuted}$, we first remove it from the mutation set $E_{mutated}$ and the selected set $E_{selected}$. Then we replace $e_{r}$ with another mutated variant $e$ that comes from the same family by a heuristic selection strategy (Section III-C2) from line 22 to line 23. Here the family refers to a set of mutated specifications $E_{f}$ that come from the same template specification. Finally, we add $e$ to the selected set $E_{selected}$ to prepare for the next iteration of verification. The above process will be repeated until all candidates in $E_{selected}$ are successfully verified i.e., $E_{refuted}$ is empty (line 16). Note that if all candidates are refuted, the process will finally select an empty set of candidates, guaranteeing the termination of the whole process.
Input : Set of template specification $E_{t}$, set of mutations $M$
Output : Set of verified specifications $E$
1 Function MutationBasedGen($E_{t},M$)
2       $E_{mutated}=\mathrm{SpecMutation}(E_{t},M)$
3       $E=\mathrm{SpecSelection}(E_{mutated},E_{t},M)$
4       return $E$
5
6Function SpecMutation($E_{t},M$)
7       $E_{mutated}=\emptyset$
8       for ${e}\in{E_{t}}$ do
9             $E_{mutated}=E_{mutated}\cup\mathrm{Mutate}(e,M)$
10
11      return $E_{mutated}$
12
13Function SpecSelection($E_{mutated},E_{t},M$)
14       $E_{selected}=E_{t}$
15       $E_{refuted}=\emptyset$
16       repeat
17             $E_{refuted}=\mathrm{Verify}(E_{selected})$
18             $E_{selected}=\mathrm{ReSelect}(E_{selected},E_{mutated},E_{refuted},M)$
19
20      until $E_{refuted}$ is $\emptyset$;
21      return $E_{selected}$
22
23Function ReSelect($E_{selected},E_{mutated},E_{refuted},M$)
24       for $e_{r}\in E_{refuted}$ do
25             $E_{mutated}=E_{mutated}\backslash\{e_{r}\}$
26             $E_{selected}=E_{selected}\backslash\{e_{r}\}$
27             $E_{f}=\mathrm{GetFamilyOf}(e_{r},E_{mutated})$
28             $e=\mathrm{SelectByHeuristic}(E_{f},M)$
29             $E_{selected}=E_{selected}\cup\{e\}$
30
31      return $E_{selected}$
32
Algorithm 1 Mutation-based Specification Generation
III-C1 Template Specification Mutation
As shown in Table I, we define four kinds of mutation operators including predicative, logical, comparative, and arithmetic. Each type of mutation corresponds to one type of operator supported by JML. LLMs perform well in capturing the overall semantics of a program, but they often make mistakes in grasping the fine-grained relationships between variables, resulting in incorrect operators used to describe the relationships between variables, which is why our mutation design is centered around the operators. When applying the mutation operator to a specification, it substitutes the operators of the corresponding type in the specification with another operator of the same type. For example, after applying a predicative mutation, a $\exists$ predicate within a specification may be substituted with $\forall$. Note that the mutation for a certain type of operator does not necessarily create only one mutated candidate. For example, the expression a <= b may be mutated to a < b or a - 1 <= b. If multiple types of mutations can be applied to a specification at the same time, we try to exhaust each combination of different types of mutations to get all potential variants. For instance, the expression x < n + 1 can be mutated to x <= n + 1 from the comparative type, x < n - 1 from the arithmetic type, or x <= n - 1 by combining them.
TABLE I: The defined mutation operators.
Mutation Type Original Operator Mutated Operators
Predicative $\forall$ $\exists$
$\exists$ $\forall$
[…]
Comparative <= <, ”- 1 <=”
>= >, ”+ 1 >=”
Arithmetic + -
III-C2 Mutated Specification Selection
Typically, for a program, Houdini [7] usually verifies all generated specifications at one time. However, a similar practice cannot be applied in SpecGen as we try to collect all potential combinations of mutations for a template specification. The set of the obtained specifications for verification is considerably large, which poses a much greater burden for the verifier within a single verification process. To address this challenge, we innovate a heuristic selection strategy to improve the stability and efficiency of verification.
In general, the heuristic selection algorithm finds a specification $\hat{e}$ such that
$\hat{e}=\mathop{\arg\max}\limits_{e\in E_{f}}\sum\limits_{m\in M}(\mathop{times}(m,e,e_{t})\cdot\mathop{weight}(m))$ (1)
where $E_{f}$ denotes a family of mutated specifications that come from the same template specification $e_{t}$, and $M$ denotes the set of all mutations. Given $E_{f}$, $e_{t}$, and $M$, we design the heuristic selection logic to prioritize selecting important candidates for verification. In particular, we assign scores for each mutated candidate $e\in E_{f}$ and select the candidate with the highest score as the output. To calculate the score of a candidate $e$, for all types of mutations $m\in M$, we sum up all the values of $times(m,e,e_{t})$ multiplied by $\mathop{weight}(m)$, where $times(m,e,e_{t})$ calculates how many times the mutation $m$ is performed when $e_{t}$ mutates into $e$, and $weight(m)$ denotes the corresponding weight of $m$.
IV Experimental Setup
We design the following four research questions for evaluation:
  • RQ1: How does SpecGen compare with the baseline approaches?
  • RQ2: How does each type of mutation contribute to the effectiveness of SpecGen?
  • RQ3: How do different candidate selection strategies affect the efficiency of SpecGen?
  • RQ4: To what extent can the generated specification contain the semantic information of the input program?
IV-A Implementation
We use the API provided by OpenAI [35] to communicate with the large language model of gpt-3.5-turbo-1106 for the experiments. We set the sampling temperature to 0.4 to balance the diversity and rigorousness of the outputs of GPT. 4 few-shot examples are used during the prompt construction to balance the input length and response time. The maximum number of rounds of conversation is set to 10. The verifier is OpenJML [21], one of the most recent JML specification verification tools, which checks the consistency between the given Java source code and the JML specifications within. Due to the incompleteness in the implementations of OpenJML, we set a timeout limit of 30 minutes for a single verification in our implementation to avoid unexpected situations, such as the non-responding of OpenJML. All experiments are conducted on an 8-core workstation with Intel Core i7-12700H CPU @2.30GHz and 32GB RAM, running Ubuntu 22.04.3 LTS. The version of OpenJDK is 1.8.0_371 for all experiments except for Houdini, which has to run under OpenJDK 1.6.0_45. We set the weight of comparative, logical, arithmetic, and predicative mutation to -1, -2, -4, and -4 respectively as the comparative mutation is more likely to pass the verification followed by the logical mutation. The predicative and arithmetic mutations are the least important through our observations from extensive experiments. Note that the weights are defined with negative values, leading to negative calculated scores as well. The reason for the design is to prioritize the specification candidates with fewer mutation occurrences.
IV-B Dataset
To comprehensively evaluate the effectiveness of SpecGen, following previous work [36], we first use an established dataset namely SV-COMP [16] benchmark for evaluation. Specifically, we used 265 class definitions in the Java category of SV-COMP benchmark and conducted the necessary modifications on part of these programs (referred to as Dataset SV-COMP hereinafter) for ease of evaluation. The remaining data in the benchmark cannot be applied for specification generation even with our modification. In particular, the verification with expected verification result FALSE (i.e., a counterexample to safety) has to be modified so that the program can exit properly, and those library calls specific to the competition settings (e.g. Verifier.nondetInt()) in the definitions are replaced with equivalent Java library calls so that they can be successfully compiled. However, after a deep analysis of the characteristics of the data from SV-COMP, we find that 88.7% programs are loop-free, indicating that the samples with more complex program structures cannot be covered by this dataset, inducing limitations on the evaluation. Also, very few datasets have been established specifically for specification generation tasks so far.
To remedy this gap, we further collect another dataset containing 120 samples as a supplement where 20 programs (including the corresponding specifications) from the dataset constructed by Nilizadeh et al. [37] and 100 programs from LeetCode [38]. This dataset is namely SpecGenBench and we have made it public [17]. The selected programs are assured of the feasibility of expressing their behaviors as JML-specified verifiable specifications. These programs involve a variety of control flow structures and encompass multiple data structures such as arrays, strings, and other data structures supported by the Java library. They also cover a diverse set of specifications including post-conditions and loop invariants, involving both linear and nonlinear relationships between variables, making them representative of a broad spectrum of scenarios. They can be categorized into five categories according to their types of control flow structures [39]. Specifically, Sequential denotes the programs without branches or loops. Branched represents the loop-free programs that will contain branches like if-else or switch structures. Single-path Loop contains the simplest type of loop, with only one layer of loop structure without branches in their loop bodies. In contrast, Multi-path Loop denotes the loops that have branches in the loop bodies. Lastly, Nested Loop denotes the programs with multiple layers of loop structure where each layer may have a branch. The quantity for Sequential, Branched, Single-path Loop, Multi-path Loop and Nested Loop is 26, 23, 24, 26, and 21, respectively.
To obtain the oracle specifications for the 100 Java programs from LeetCode, we follow a similar procedure in Nilizadeh et al. [37] with the help of human experts. Three experts with rich experience in formal verification were employed, to manually write specifications for each program. Each expert is required to write the specifications that can be successfully verified to describe the functionality and behavior of the program as accurately and comprehensively as possible. For a single program, if multiple experts have written verifiable specifications, another expert is responsible for selecting one of them as the oracle.
As Daikon [9] requires a set of test suites instrumented into the source code to execute the code, we also manually write these test suites for it. Lastly, we instrument the dummy function calls at the top of each loop body in a program to ensure Houdini and Daikon can generate the specifications at these program points.
IV-C Baselines
We select two conventional approaches and a purely LLM-based approach as the baselines for comparison.
Houdini [7]. It is a template-based JML annotation generator that relies on a series of pre-defined templates to generate candidate specifications. Given the input program, it first generates candidate specifications by filling in the templates with available variables and all kinds of operators. Afterward, It iteratively invokes a JML specification verifier to check their correctness and removes the refuted ones. The process will be repeated until the remaining candidates are all verified.
Daikon [9]. It is a classic tool for the dynamic detection of program specifications which relies on the dynamic execution trace of the target program to infer likely specifications. Given the input program, it first instruments the target program to trace certain variables and extracts execution traces. Then the inference engine reads the trace data and infers the potential invariants with a generate-and-check algorithm. Daikon supports dynamic detection for Java, C/C++, C#, and Perl programs along with various formats such as DBC format [40], JML format [18], and CSharpContract format [41].
Apart from the above approaches used for specification generation, we further add the purely LLM-based approaches.
Purely LLM-based Approaches. They refer to those relying entirely on the large language model i.e., gpt-3.5 in our work to generate specifications (Section III-B). We compare SpecGen with few-shot setting and conversational setting. We set 0-shot, 2-shots, and 4-shots for few-shot comparison. The conversational setting is also based on 4-shot examples and the error message is provided for conversation. The other settings are the same with SpecGen.
IV-D Evaluation Metrics
Following the previous works [36, 42], we use the metric of Number of Passes for assessment. We further add more metrics for comprehensive evaluation.
Number of Passes. It defines the number of programs for which the generated specifications of an approach pass the validation by the verifier. For a program, we consider the specifications that pass the verifier as the correct specifications.
Success Probability. It is used to evaluate the model-based approaches. The randomness inherent in the content generated by large language models may introduce a certain level of contingency in a successful generation. Thus, we use the success probability for the measurement. For a test program, it is calculated by $\frac{N_{sucess}}{N_{attempt}}$ where $N_{sucess}$ denotes the number of successful generations of verifiable specifications and $N_{attemp}$ denotes a fixed number of trials in total (10 times in SpecGen).
Number of Verifier Calls. It is used to evaluate the efficiency of our approach. We propose a heuristic selection algorithm to prioritize the important candidates for the verifier to verify. To evaluate the effectiveness of the proposed selection algorithm, we use the number of verifier calls as the evaluation metric.
User Rating. It aims to measure the semantic quality of the generated specifications. We invited 15 Ph.D. students who are experts in Java programming language to rate the specifications generated by different approaches. The research is conducted using the Likert Scale [43], where students are required to give a rating from one point to five points for each case. A reference rating criteria (can be accessed at [17]) is also presented to the students.
V Experimental Results
V-A RQ1: Comparison with Baselines
TABLE II: Number of programs that successfully pass the verifier and average success probability.
Approach SV-COMP (265) SpecGenBench Overall (385)
Sequential
Branched
Single-path Loop
Multi-path Loop
Nested Loop
Num. Prob. Num. Prob. Num. Prob. Num. Prob. Num. Prob. Num. Prob. Num. Prob.
Daikon 51 - 10 - 10 - 0 - 1 - 0 - 72 -
Houdini 56 - 14 - 11 - 10 - 4 - 3 - 98 -
Purely LLM-based 0-shot 81 18.28% 23 74.19% 17 $5$ 7.08% 7 18.13% 2 3.33% 135 22.93%
2-shot 83 18.79% 20 61.06% 17 53.91% 8 19.29% 13 26.58% 4 3.81% 145 23.48%
4-shot 94 19.40% 23 73.85% 20 57.33% 10 23.40% 12 24.95% 5 6.24% 164 25.25%
Conversational 146 30.95% 23 82.49% 20 75.43% 12 27.02% 13 35.38% 4 9.20% 218 35.95%
SpecGen 179 40.41% 24 92.31% 20 79.57% 23 73.75% 20 60.38% 13 36.55% 279 59.97%
The experimental results are presented in Table II where Num. denotes the number of successfully handled programs, and Prob. denotes the average success probability. For the purely LLM-based approaches and SpecGen, the number of trials for each program is 10 times.
Performance on SV-COMP. From Table II, we can find that on the SV-COMP dataset (265 programs in total), Houdini handled 56 programs, which is more than Daikon. However, both of them underperform LLM-based approaches even in LLM’s simplest setting i.e. 0-shot setting, which handled 81 programs, demonstrating the feasibility of employing LLMs for formal program specification generation. Among the LLM-based approaches with few-shot settings, as the number of given few-shot examples increases, the number of programs that LLM can generate verifiable specifications is also increased, substantiating the effectiveness of the few-shot examples. Based on the 4-shots examples in the initial prompt, the multi-turn conversational manner (in Section III-B) can further improve the performance, with 146 programs handled, compared to 94 programs of 4-shot LLM. Lastly, SpecGen outperforms the conversational-based approach, with the number of programs that generated verifiable specifications increasing from 146 to 179.
Performance on SpecGenBench. For deeper insights into the performance of different approaches on different types of programs, we further investigate the results on SpecGenBench. Similar to SV-COMP, Houdini outperforms Daikon, especially in generating specifications for complex program structures such as loops, Houdini exhibits certain abilities. In terms of the LLM-based approaches, we find that both SpecGen and the conversational-based approach are competent in generating specifications for relatively simple programs such as sequential and branched. However, when it comes to generating more complex program structure specifications such as loop, SpecGen has a clear advantage. The benefits are from our designed mutation-based specification generation (Section III-C), which can correct the erroneous output of the large language model to generate the verifiable specifications. Although SpecGen can generate more accurate specifications for different loop structures, it has a relatively poor performance in generating verifiable specifications that the program has the nested loop.
For a program, we further use the metric namely success probability to evaluate the probability of successful generation for this program by LLMs in 10 times trials due to the randomness inherent in the generated content by LLMs. We can observe that SpecGen achieves an overall success probability of 59.97%, which is significantly higher than the value of conversational manner i.e., 35.95%. It indicates that compared with the purely conversational manner, SpecGen is more reliable with higher probabilities and lower chances of randomness to generate verifiable specifications. From the results from SV-COMP and SpecGenBench, we can find that our proposed approach is orthogonal to different datasets.
Figure 5: Venn diagram of verifiable programs.
To the end, the Venn diagram of the programs in SpecGenBench for which SpecGen and baseline approaches successfully generated verifiable specifications is presented in Fig. 5. We merge the programs successfully handled by different purely LLM-based approaches together as Purely LLM-based in Fig. 5. It is noteworthy that SpecGen generates verifiable specifications for 14 programs that other baselines fail to generate, where 6 are from the Nested Loop, 3 are from the Multi-path Loop, and 5 are from the Single-path Loop. Further investigation of these 14 programs, we find that they are relatively complicated, making it hard for the baseline approaches to correctly generate. The generated specifications of different approaches for these programs are presented on our website [17]. We also present a case study to illustrate the successful/unsuccessful generation on the website.
RQ1: SpecGen significantly outperforms the baselines on two different datasets, generating verifiable specifications for 279 out of 385 programs, whereas the conversation-driven approach, Houdini and Daikon succeeds in 218, 98, and 72 respectively.
V-B RQ2: Ablation Study on Mutation Types
We conduct an ablation study to evaluate the effectiveness of different mutation types in SpecGen. The results are shown in Table III where w/o {*} denotes the disabled mutation type.
We can find that SpecGen successfully generates verifiable specifications for 279 out of a total of 385 programs. SpecGen w/o Comparative addresses the least number of programs i.e., 223, indicating that the comparative mutation is the most important in the defined mutation operations. The main reason is the frequent usage of numerical variables in programs and the recurring need to bound their range in the specifications. SpecGen w/o Logical has the second least number of programs i.e., 235, indicating that the logical operators are also important to generate verifiable specifications. This is due to the necessity of combining two or more expressions for different properties with logical operations when specifying complex behaviors. SpecGen w/o Predicative and SpecGen w/o Arithmetic have the most number of programs (259 and 262 respectively), which means both of them are less important than the comparative and logical mutation. We still consider them as they are applicable in certain situations. In some complex programs, specifications generated by the LLM are prone to have predicate errors. Hence, the predicative mutation will be useful. Similar cases exist when there are complicated numerical constraints on variables, where mutations on arithmetic operators turn out to be helpful.
Further analyzing the effectiveness of the mutation type for different kinds of programs, we can find that SpecGen w/o Comparative handles an especially lower number of programs in the loop category including single-path, multi-path, and nested loop. It is due to the rigid demand of scope bounding for loop variables when loops are involved. Scope bounding for loop variables is an intricate work where LLMs frequently make mistakes, substantiating the importance of comparative mutations. The performance of SpecGen w/o Predicative also drops on programs with nested loop, because of the relatively higher quantity and complexity of $\forall$ and $\exists$ statements involved in these nested programs.
TABLE III: Effectiveness of different types of mutations.
Approach SpecGenBench SV-COMP (265) Total (385)
Sequential
Branched
Single-path
Multi-path
Nested
w/o Predicative
24 20 20 19 9 167 259
w/o Logical
24 18 14 18 10 151 235
w/o Comparative
24 19 13 12 7 148 223
w/o Arithmetic
23 19 18 21 11 170 262
SpecGen 24 20 23 20 13 179 279
RQ2: Each type of mutations contributes differently to SpecGen. The comparative mutation contributes the most to the performance while the predicative and arithmetic are less important. When combining them together, SpecGen achieves the best performance.
TABLE IV: Average numbers of verifier calls in a single run under different specification selection strategies in Section III-C2.
Strategy SV-COMP SpecGenBench Total
Sequential Branched
Single-path
Multi-path
Nested
Random 9.41 2.70 2.32 32.62 55.16 41.97 18.44
Heuristic 8.91 2.59 2.16 24.21 43.58 34.99 15.51
V-C RQ3: Effectiveness of Selection Strategy
In Section III-C2, we design the heuristic selection strategy to improve the efficiency of verification. We also conduct an experiment to compare with the random selection strategy. Specifically, when a candidate specification is refuted by the verifier, we randomly select another specification from $E_{f}$ for replacement. To compare with different strategies, for each program that successfully generates verifiable specifications by SpecGen, we run SpecGen 5 times to obtain the average number of verifier calls as the evaluation metric.
Performance on SV-COMP. Using the random selection strategy makes 9.41 verifier calls on average while the heuristic selection strategy takes 8.91 calls, resulting in an improvement of 5.30%. The improvement is relatively modest and the values denote that only less than 10 verifier calls on average are used to generate verifiable specifications for programs in SV-COMP. The main reason is that the number of rounds for conversation is set to 10 in SpecGen, the specifications for these programs tend to be successfully generated within the conversation module (Section III-B). It is before the selection strategy used in the mutation-based specification generation (Section III-C) comes into effect.
Performance on SpecGenBench. Using the random selection strategy takes 36.20 verifier calls on average while the heuristic selection strategy takes 28.51 calls in terms of five categories in SpecGenBench. Hence, the heuristic selection strategy achieves an improvement of 21.23%. Furthermore, we can observe that the improvements in different categories of programs vary significantly. The improvements in the loop categories including single-path, multi-path, and nested loop are more significant than sequential and branched. The main reason is that generating specifications for loop-containing programs is challenging, and usually requires more iterations to obtain verifiable specifications. In this case, a good selection strategy often highlights advantages more effectively. However, the improvements in sequential and branched categories are fewer. The reason is similar to SV-COMP. The high efficiency of conversational generation on sequential and branched programs makes the selection strategies invalid. Nevertheless, loop structures are common in programs, thus a heuristic selection strategy to improve the validation efficiency is still helpful and necessary.
RQ3: The heuristic selection strategy effectively improves the efficiency of SpecGen. It is especially useful to generate the specifications for programs that have more complex structures such as loops.
V-D RQ4: User Study on the Quality of Specifications
To evaluate the semantic quality of the generated specifications, we conduct a user study for evaluation. 15 Ph.D. students are invited to rate the specifications generated by different approaches. A detailed description of the questionnaire and rating criteria is given in Section IV-D. We selected the 15 programs from the dataset of SpecGenBench that can be handled by all of Houdini, Daikon, and SpecGen. Apart from the specifications generated by these approaches, we also add the ground truth as a reference. The rating scores are presented in Table V, where the average score for full marks is 5.
We can observe that the ground truth specifications (oracle) receive an average rating score of 4.83, indicating that the semantics of these programs can be described comprehensively through JML specifications. Furthermore, the specifications generated by SpecGen received a rating score of 4.54, which is close to the oracle, indicating that the generated specifications by SpecGen can also describe the real behaviors of the input program more fully. Among the 15 programs, all rating scores given to SpecGen are above 3, with the lowest rating being 3.33, meaning that in the worst case, SpecGen can still generate non-trivial specifications about the properties of the input program. Benefiting from the code comprehension capability of LLMs, SpecGen can generate specifications closely related to the real behaviors of the input program. In comparison, the specifications generated by Houdini and Daikon received an average rating score of 2.32, reflecting the semantic weakness in these specifications, which is consistent with the claims in Section II-B. As Houdini and Daikon rely on pre-defined templates for generation, the specifications produced by them are often simplistic and trivial, making it challenging to precisely capture the actual behavior and functionality of the input program. We further provide a rating of 100 programs that SpecGen can successfully generate verifiable specifications following the criteria from Section IV-D on our website [17].
RQ4: SpecGen received an average rating score of 4.54, which is close to the 4.83 of the oracle specifications, demonstrating the ability of SpecGen to accurately characterize the real behaviors of the input program and generate specifications with comprehensive program semantics.
TABLE V: Average rating scores on the generated specifications by different approaches.
Test case Houdini Daikon SpecGen Oracle
Absolute 3.50 3.36 4.85 5.00
AddLoop 2.40 1.33 4.57 5.00
Conjunction 4.50 3.50 5.00 5.00
ConvertTemperature 2.33 2.50 5.00 5.00
Disjunction 2.50 3.50 5.00 5.00
FizzBuzz 2.63 2.86 5.00 5.00
IsCommonFactor 2.00 4.13 4.14 4.71
IsPalindrome 1.83 1.17 4.75 5.00
IsSubsequence 2.43 1.13 4.14 4.00
IsSuffix 2.20 1.50 4.33 4.63
MulLoop 1.88 1.25 3.33 5.00
MySqrt 2.00 2.80 3.75 4.25
Perimeter 1.00 2.80 4.78 5.00
SmallestEvenMul 2.57 1.00 4.50 5.00
Swap 1.00 2.00 5.00 4.88
Average 2.32 2.32 4.54 4.83
VI Threats to Validity
Internal Validity. First, the prompts we used to communicate with the LLM may affect our results. To mitigate it, we refer to Xia et al. [11] to design the prompt. We plan to investigate the effect of different prompts in the future. Second, a potential threat lies in the risk of data leakage. Our constructed dataset SpecGenBench consists of 100 programs with expert-written specifications and 20 programs with their corresponding specifications from Nilizadeh et al. [37]. The former does not have the issue of data leakage as the specifications are written by experts in our research. However, since gpt-3.5 does not release its model as well as the training data, the latter 20 programs from the existing dataset may have the risk. The used dataset SV-COMP also has this risk. Nevertheless, through our observation of SpecGen on these programs, we have never spotted a situation where the output of SpecGen is the same as the existing oracle. Hence, we believe this threat is limited. Furthermore, even if we remove these potentially risky programs, SpecGen still successfully handles 87 programs in the remaining 100 programs, which is also the best.
External Validity. One of the external threats lies in the accuracy of the verifier (OpenJML). Due to the implementation flaws in OpenJML, there may be cases where some correct specifications fail to pass verification. This is an inevitable problem that other verifiers [44] have to face as well. The reason lies in the undecidability of automatic software verification [45, 46]. Even in such a situation, SpecGen achieves impressive performance with the majority of testcases successfully generated verifiable specifications. Another threat is the potential bias of the hand-written specifications by experts. To mitigate this, we follow the procedure in Nilizadeh et al. [37]. First, the selected experts should have rich experience in writing specifications. Second, the initially chosen specifications should be verifiable by the verifier. Last, if multiple experts have written the specifications that pass the verifier, another expert is responsible for selecting one.
VII Related Work
Large Language Models. With the advancement of generative AI, Large Language Models (LLMs) have emerged as a formidable force and have quickly found widespread applications. LLMs are characterized by their immense parameter scale and training dataset size [47]. An important feature of LLMs is their ability for in-context learning [32], which enhances the coherence between the context and the output of LLMs. The learning ability gives rise to a unique usage of LLMs known as prompting [48], where a natural language description of the intended downstream task is provided to the LLM before assigning it the task. LLMs initially demonstrated remarkable capabilities in the field of Natural Language Processing (NLP) [49], excelling in tasks such as document classification [50], text summarization [51], and machine translation [52]. They are also widely deployed in various software engineering tasks [14, 10], including software testing [25, 26], code generation [13, 28] and code summarization [12]. Compared with these works, our goal is to employ LLMs for the automated generation of program specifications, which is important in formal methods.
Program Specification Generation. The research on program specification generation can be categorized into two types: natural language specification generation and formal specification generation. Natural language specification generation primarily manifests as code summarization [53], a process of automatically generating accurate, human-readable descriptions of code functionality. Numerous efforts have been made to utilize machine learning methods for code summarization [54, 55, 56, 57]. Formal specification generation primarily takes the form of the generation of program invariants, the formal language representations of properties that a program is guaranteed to satisfy at a certain program point. In invariant generation, a large amount of research focuses on the generation of loop invariants [58, 59, 60, 61, 62, 63, 64], while the rest of the works attempt to generate invariants of other forms, e.g. pre-conditions [65, 42], post-conditions [66, 67, 36, 68, 69], assertion-based invariants [70] and finite automata [71, 72, 73]. There also have been efforts employing fine-tuning methods to help LLMs better adapt to specification generation [74]. Among these works, artifacts of Ghosal et al. [42] and Pei et al. [74] are not publicly available, the artifact of Alshnakat et al. [36] is built for C code and Frama-C contracts, and the grammar for specifications of Molina et al. [68, 69] involves specific features, which cannot be trivially translated into equivalent JML, thus we cannot include them for comparison. SpecGen utilizes the code comprehension capability of LLMs for program specification generation in a conversational manner, further followed by the mutation-based approach for enhancement.
VIII Conclusion
In this paper, we introduced SpecGen, a novel approach that utilizes the Large Language Model for formal program specification generation. Leveraging the code comprehension ability of LLMs as well as the well-designed mutation-based specification generation component, our approach is capable of accurately capturing the behaviour and functionality of input programs to generate accurate specifications. A comprehensive evaluation between SpecGen and other baselines is conducted on two different datasets, the benchmark for the Java category of SV-COMP, and a more diverse and manually constructed dataset containing 120 programs. The extensive experimental results haver demonstrated that our approach significantly outperforms the baseline approaches.